General Announcement: - Harald Sinn inform on the first lasing for XFELO on Sunday morning shift; 100 bunches in the pulse train was giving synchrotron already on Saturday showing the cavity was aligned. On Sunday was difficult to find FEL intensity but successful by ~14:30 while scanning cavity. After getting the signal, it was surprisingly stable and running well. CTRL: - Nothing major to report. DA: - Nothing to report. EEE: - Nothing to report. VAC: - Nothing to report. XPD: - No issues from XPD; - Success with FD activities. DET: - Nothing to report. ITDM: - Nothing to report. XRO: - Not present. LAS: - Not present. UNSYS: - Yesterday a ZZ was done in XTD4 to fix SASE3 cell26 Phase shifter. SPB: - Issue with Laser interlock in laser hutch; it will be discussed with Pouneh S. soon. FXE: - Not present. MID: - All fine at MID. HED: - Nothing to report. SCS: - Not present. SQS: - All good. SXP: - Nothing to report.